# Divergence-Based Similarity Function for Multi-View Contrastive Learning

### v1

@@ 4.4 Divergence-based similarity function
4.4.1 Similarity function
Similarity function $\text{sim}(\cdot,\cdot)$ is given by $\text{sim}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. For example, cosine similarity function is given as $\text{sim}_{\text{cos}}(\cdot,\cdot):\mathbb{S}_{p-1}\times\mathbb{S}_{p-1}\rightarrow\mathbb{R}$, $\text{sim}_{\text{cos}}(x,y)=x^{T}y$. Accordingly, DSF handles multi-view features, and its similarity function is defined as $\text{sim}_{\text{div}}(\cdot,\cdot):\mathbb{S}_{p-1}^{m}\times\mathbb{S}_{p-1}^{m}\rightarrow\mathbb{R}$.
4.4.2 KL Divergence of vMF Distribution
Let $\mathcal{D}_{i}=\mathcal{D}(\mu_{i},\kappa_{i})$ and $\mathcal{D}_{j}=\mathcal{D}(\mu_{j},\kappa_{j})$ be any two vMF distribution. The KL divergence value of $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ is expressed as the sum of three terms.
$\begin{split}&D_{\mathrm{KL}}(\mathcal{D}_{i}||\mathcal{D}_{j})\\ &=\log\frac{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{i},\kappa_{i})}{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{j},\kappa_{j})}\\ &=\left(\frac{p}{2}-1\right)\log\frac{\kappa_{i}}{\kappa_{j}}+\log\frac{I_{\frac{p}{2}-1}(\kappa_{j})}{I_{\frac{p}{2}-1}(\kappa_{i})}\\ &\qquad\qquad\qquad+A_{p}(\kappa_{i})\cdot\left(\kappa_{i}-\kappa_{j}\cdot\mu_{i}^{T}\mu_{j}\right)\\ \\ \end{split}$ (9)
The KL divergence range in from 0 to infinity. KL divergence is not a distance function of two distribution. As the KL divergence value close to zero, it means the distributions are similar i.e. $\mu,\kappa$ become close. So, our proposed divergence-based similarity function is negative KL divergence value as following.
@@ 4.4.2 KL Divergence of vMF Distribution
Here, $\mathcal{D}_{i}^{+}$ denotes the positive distribution paired with $\mathcal{D}_{i}$. The negative distributions are denoted by $\mathcal{D}_{j}$, where $K$ represents the number of negative distributions. Note that the value of $K$ may vary depending on the training framework.
@@ 4.5 Analysis : Relation with Cosine Similarity Function
Next, we establish the connection between the conventional cosine similarity function and our proposed DSF. We further explain why cosine similarity requires a temperature scaling term for effective optimization, whereas our DSF inherently eliminates the need for such adjustment.
@@
Let $Z_{i}=\{z^{1}_{i},\cdots,z^{m}_{i}\}$ and $Z_{j}=\{z^{1}_{j},\cdots,z^{m}_{j}\}$ are $m$ multi-view features. $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ are the estimated von Mises-Fisher distribution from $Z_{i}$ and $Z_{j}$. Let $\text{sim}_{\text{cos}}(\cdot,\cdot)$ is cosine similarity and $\text{sim}_{\text{div}}(\cdot,\cdot)$ is the negative Kullback-Leibler divergence between two vMF distribution. If m=1, $\kappa_{1}$=$\kappa_{2}$ and $A_{p}(\kappa_{1})\cdot\kappa_{1}=1/\tau$ then
$\text{sim}_{\text{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\text{sim}_{\text{cos}}(z_{i},z_{j})/\tau+constant$ (12)
Furthermore, InfoNCE Loss with $sim_{\text{div}}$ becomes equivalent to that using $sim_{\text{cos}}$ one:
$L^{\text{div}}_{\text{InfoNCE}}=L^{\text{cos}}_{\text{InfoNCE}}$ (13)
@@
Proof is at Appendix A. ∎
Theorem 1 says that when we have two views and kappa is set under certain conditions. Then our proposed divergence-based simiarity function becomes cosine simiarity function with some constant value. Furthermore, this leads to fact that InfoNCE loss becomes equivalent.
Now, we explain the dependency of cosine similarity on temperature hyperparameter and independence of our DSF with respect to temperature hyperparameter.
@@
Let $K$ be the number of negative pairs, and assume all negative similarities have the same value. Let $s^{+}$ and $s^{-}$ denote the positive and negative similarities, respectively. Then, the InfoNCE loss becomes:
$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(s^{+}-s^{-})}{\exp(s^{+}-s^{-})+K}\\ (s^{+}-s^{-})&\in\left\{\begin{matrix}[-\frac{2}{\tau},\frac{2}{\tau}]&\text{if }\text{sim}=\text{sim}_{\text{cos}}\\ (-\infty,\infty)&\text{if }\text{sim}=\text{sim}_{\text{div}}\end{matrix}\right.\end{split}$ (14)
@@
$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(2/\tau)}{\exp(2/\tau)+K}\\ &=6.3196,\quad\tau=1.0,\,K=4096\\ &=4.3310,\quad\tau=0.5,\,K=4096\\ &=0.1706,\quad\tau=0.2,\,K=4096\\ &=0.0000,\quad\tau=0.1,\,K=4096\\ \end{split}$ (15)
It has been theoretically shown (see Proposition1) that cosine similarity cannot represent a appropriate loss value without the use of a temperature term $\tau$. The core issue arises from the limited range of the similarity margin $(s^{+}-s^{-})$, which restricts the contrast between positive and negative pairs. To compensate for this limitation, the temperature parameter $\tau$ is introduced to amplify this margin, allowing for better gradient signals during training.
The number of negative samples $K$ varies depending on the contrastive learning framework. For instance, in SimCLR and DCL, $K=2B-2=254$ when the batch size $B=128$, while in MoCo, $K$ is determined by the size of the memory queue, ranging from 4,096 to 65,536 depending on the dataset.
As summarized in Table 2, the InfoNCE loss yields inappropriate values without the temperature parameter $\tau$ in all $K$ values. It says cosine similarity requires a carefully tuned temperature parameter across all frameworks to function effectively. In contrast, our proposed DSF expresses an appropriate loss without the need for such a temperature hyperparameter. See Figure 5 for $(s^{+}-s^{-})$ margin in our DSF experiment.
@@
Table 2: InfoNCE loss values under optimal similarity ($s^{+}=1$, $s^{-}=-1$) for varying temperature $\tau$ and the number of negative pairs $K$.
@@
Input: Instance $x_{i}\in\mathbb{R}^{C\times H\times W}$, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$.
@@
2:  Extract features:
$z_{i}^{l}=\frac{f_{\theta}(x_{i}^{l})}{\|f_{\theta}(x_{i}^{l})\|},\quad z_{B+i}^{l}=\frac{f_{\theta^{\prime}}(x_{B+i}^{l})}{\|f_{\theta^{\prime}}(x_{B+i}^{l})\|}\quad$
for $\text{for }l=1,\dots,m$
3:  if multi-view method = loss_avg:
$L_{\text{loss\_avg}}=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}L_{\text{InfoNCE}}^{\cos}(z_{i}^{l},z_{B+i}^{l^{\prime}})$
4:  if multi-view method = fea_avg:
$L_{\text{fea\_avg}}=L_{\text{InfoNCE}}^{\cos}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$
5:  if multi-view method = DSF:  Estimate vMF distribution parameters as in Eq. (3)–(4):
$(\mu_{i},\kappa_{i})\leftarrow\{z_{i}^{1},\dots,z_{i}^{m}\}$
$(\mu_{B+i},\kappa_{B+i})\leftarrow\{z_{B+i}^{1},\dots,z_{B+i}^{m}\}$
6:   Compute divergence-based similarity:
$\mathrm{sim}_{\mathrm{div}}(\mathcal{D}_{i},\mathcal{D}_{B+i})=-D_{\mathrm{KL}}(\mathcal{D}_{i}\,\|\,\mathcal{D}_{B+i})$
$L_{\text{DSF}}=L_{\text{InfoNCE}}^{\mathrm{div}}\left(\mathcal{D}(\mu_{i},\kappa_{i}),\mathcal{D}(\mu_{B+i},\kappa_{B+i})\right)$
@@ 5.1 Experimental Setup
Dataset
This study utilizes two benchmark datasets: CIFAR-10 and ImageNet-100. CIFAR-10 [Krizhevsky et al., 2009] consists of 60,000 color images of size $32\times 32$ pixels, divided into 50,000 training and 10,000 test images across 10 object classes. ImageNet-100 is a curated subset of the ImageNet dataset, comprising 100 randomly selected classes with approximately 1,300 images per class. It serves as a mid-scale benchmark for evaluating representation learning performance.
These datasets are used to validate the proposed method on both small-scale and mid-scale visual recognition tasks.
Model Arichitecture
We primarily follow the MoCo v2 [Chen et al., 2020b] framework for contrastive learning. Throughout this paper, any reference to ”MoCo” refers to MoCo v2 unless otherwise specified. We adopt ResNet-18 as the encoder for experiments on CIFAR-10 The final fully connected layer projects features into a 128-dimensional embedding space.
For large-scale evaluation, we extend our experiments to ImageNet-100 using MoCo v3 [Chen et al., 2021] with a Vision Transformer (ViT-S/16) backbone. This enables assessment of the scalability of the proposed method across both convolutional and transformer-based architectures.
Data Augmentation
For data augmentation, we apply a combination of RandAugment [Cubuk et al., 2020], ColorJitter, Random Grayscale, Gaussian Blur, and Horizontal Flip, following common practice in contrastive learning setups.
@@ 5.2 Evaluations
To evaluate the model’s performance, we conducted following tasks: k-Nearest Neighbors (kNN) and Linear Evaluation. Linear Evaluation assesses the representational quality of the encoder by attaching a classifier, while kNN measures similarity-based performance directly within the encoder’s embedding space.
@@
We evaluated the model with a kNN approach to measure the encoder’s capacity to retain similarity in the embedding space. Evaluation was conducted on CIFAR-10, ImageNet-100 using the output feature of the trained query encoder. The K value was set to 200, enabling an assessment of the encoder’s representations in a non-parametric, unsupervised manner. In our experiments, the proposed model achieved a kNN accuracy of 79.00% on the ImageNet-100 dataset, outperforming the baseline MoCo v3 model, which attained 74.08%. On the CIFAR-10 dataset, our model achieved 90.04% accuracy, compared to 88.44% with the MoCo baseline.
@@
To further validate the effectiveness of our method, we evaluate the quality of the learned representations through linear evaluation. Specifically, after performing self-supervised pretraining on ImageNet-100 and CIFAR-10 using our proposed DSF framework, we freeze the encoder and train a linear classifier (a single-layer MLP) on top of the extracted features.
On the ImageNet-100 dataset, our method achieved a linear evaluation accuracy of 85.66%, outperforming the baseline MoCo v3 model, which attained 81.52%. Similarly, on the CIFAR-10 dataset, our method achieved an accuracy of 91.21%, compared to 88.10% obtained by the MoCo baseline. These results demonstrate that DSF leads to stronger and more transferable representations across different datasets.
(a) kNN Top-1 Accuracy with respect to number of view.
(b) Linear Evaluation Top-1 Accuracy with respect to number of view.
@@ 5.3 Same Resource Comparison
Our method samples $M$ multiple augmented views from each image, which could potentially increase computational overhead. Therefore, it is essential to ensure that the proposed approach remains efficient under equivalent computational constraints.
Computational Efficiency.
Our method samples $M$ augmented views from each image instance, which could potentially increase computational overhead. To address this, we ensure fair comparisons by maintaining consistent computational budgets across all experiments.
To ensure the practicality of our method, we carefully control and match computational budgets across all experiments. Specifically, we fix the product of batch size and number of views—i.e., $B\times M$—so that memory usage remains constant regardless of the number of views. For example, in our ImageNet-100 experiments (see Table 1), the baseline MoCo is trained with a configuration of $512\times 2$, whereas our method and other multi-view baselines use $128\times 8$, maintaining identical GPU memory usage. This configuration strategy is also consistently applied to CIFAR-10 experiments.
Moreover, to equalize the overall training time, we proportionally reduce the number of training epochs for multi-view methods. As a result, the total runtime difference between our approach and baseline models is strictly limited to within 30 minutes—less than 4% and 10% of the entire training duration in the ImageNet-100 and CIFAR-10 setting. These measures ensure that the observed performance improvements are attributable to our proposed DSF method itself, rather than to increased computational resources.

### v2

@@ 4.3 Divergence-based similarity function
Similarity function
Similarity function $\text{sim}(\cdot,\cdot)$ is given by $\text{sim}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. For example, the cosine similarity function is given as $\text{sim}_{\text{cos}}(\cdot,\cdot):\mathbb{S}_{p-1}\times\mathbb{S}_{p-1}\rightarrow\mathbb{R}$, $\text{sim}_{\text{cos}}(x,y)=x^{T}y$. Accordingly, we define the DSF similarity function to handle multi-view features as $\text{sim}_{\text{div}}(\cdot,\cdot):\mathbb{S}_{p-1}^{m}\times\mathbb{S}_{p-1}^{m}\rightarrow\mathbb{R}$. More precisely,
[…]
KL Divergence of vMF Distribution
Let $\mathcal{D}_{i}=\mathcal{D}(\mu_{i},\kappa_{i})$ and $\mathcal{D}_{j}=\mathcal{D}(\mu_{j},\kappa_{j})$ be any two vMF distributions. The KL divergence value between $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ is expressed as the sum of three terms:
[…]
Figure 7 in the Appendix illustrates the behavior of each term as well as the overall KL divergence values. Detailed explanation and evaluation of $A_{p}(\cdot)$ and $I_{p}(\cdot)$ are provided in Appendix C.
The KL divergence ranges from 0 to infinity. As the KL divergence approaches zero, the two distributions become similar, indicating that $\mu_{i}\approx\mu_{j}$ and $\kappa_{i}\approx\kappa_{j}$. Conversely, a large KL divergence implies that the two distributions are highly dissimilar. Therefore, we define the DSF as the negative KL divergence. We use following expression for notational convenience throughout the remainder of the paper:
@@
Here, $\mathcal{D}_{i}^{+}$ denotes the positive distribution paired with $\mathcal{D}_{i}$. The negative distributions are denoted by $\mathcal{D}_{j}$, where $K$ represents the number of negative distributions. Note that the $K$ may vary depending on the training framework.
@@ 4.4 Analysis : Relation with Cosine Similarity Function
Next, we establish the connection between the conventional cosine similarity function and DSF.
@@
If $m=1$ and all concentration parameters $\kappa$ have the same value $\kappa^{*}$ and $A_{p}(\kappa^{*})\cdot\kappa^{*}=1/\tau$ then,
$\mathrm{sim}_{\mathrm{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\mathrm{sim}_{\mathrm{cos}}(z_{i},z_{j})/\tau+constant$ (13)
Furthermore, InfoNCE Loss with $\mathrm{sim}_{\mathrm{div}}$ becomes equivalent to that using $\mathrm{sim}_{\mathrm{cos}}$ one:
$L^{\mathrm{div}}_{\mathrm{InfoNCE}}=L^{\mathrm{cos}}_{\mathrm{InfoNCE}}$ (14)
@@
See Appendix A. ∎
Theorem 1 states that when we have two views and $\kappa$ is set under certain conditions, DSF becomes cosine similarity with some constant value. Furthermore, this leads to the fact that InfoNCE losses become equivalent.
We next provide an explanation of why cosine similarity requires a temperature scaling term for effective optimization, while DSF inherently eliminates the need for such adjustment.
@@
Let $K$ be the number of negative pairs, and assume all negative similarities have the same value. Let $s^{+}$ and $s^{-}$ denote the positive and negative similarity, respectively. Then, the InfoNCE loss becomes:
[…]
@@
[…]
We show (see Proposition 1) that cosine similarity alone fails to produce appropriate loss values without a temperature term $\tau$, due to the limited similarity margin $(s^{+}-s^{-})$ between positive and negative pairs. The temperature parameter compensates for this by amplifying the margin. As shown in Table 2, InfoNCE loss with cosine similarity produces suboptimal values for all $K$ when $\tau$ is omitted, highlighting the need for careful tuning. In contrast, DSF achieves appropriate loss values without requiring such tuning. See Figure 5 for similarity in DSF.
@@
Table 2: InfoNCE loss values under optimal similarity ($s^{+}=1$, $s^{-}=-1$) for varying temperature $\tau$ and the number of negative pairs $K$. The number of negatives $K$ depends on the framework: SimCLR uses K=2B2=254K=2B-2=254 when $B=128$, while MoCo sets $K$ by the memory queue size (4096–65536).
@@
Input: Data $x_{i}\in\mathbb{R}^{C\times H\times W}$, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$.
@@
2: Extract features: $z_{i}^{l}=\frac{f_{\theta}(x_{i}^{l})}{\|f_{\theta}(x_{i}^{l})\|},\,z_{B+i}^{l}=\frac{f_{\theta^{\prime}}(x_{B+i}^{l})}{\|f_{\theta^{\prime}}(x_{B+i}^{l})\|}\,\text{for }l=1,\dots,m$
3:  if multi-view method = Loss_avg:   $L_{i,\text{Loss\_avg}}=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}L_{i,\text{InfoNCE}}^{\cos}(z_{i}^{l},z_{B+i}^{l^{\prime}})$
4:  if multi-view method = Fea_avg:   $z_{i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l},\quad z_{B+i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{B+i}^{l}$   $L_{i,\text{Fea\_avg}}=L_{i,\text{InfoNCE}}^{\cos}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$
5:  if multi-view method = DSF:   Estimate vMF distribution parameters as in  Eq. (3)–(4):    $(\mu_{i},\kappa_{i})\leftarrow g_{\text{est}}\left(\{z_{i}^{1},\dots,z_{i}^{m}\}\right)$    $(\mu_{B+i},\kappa_{B+i})\leftarrow g_{\text{est}}\left(\{z_{B+i}^{1},\dots,z_{B+i}^{m}\}\right)$
6:   Compute divergence-based similarity:    $\mathrm{sim}_{\mathrm{div}}(\mathcal{D}_{i},\mathcal{D}_{B+i})=-D_{\mathrm{KL}}(\mathcal{D}_{i}\,\|\,\mathcal{D}_{B+i})$    $L_{i,\text{DSF}}=L_{i,\text{InfoNCE}}^{\mathrm{div}}\left(\mathcal{D}(\mu_{i},\kappa_{i}),\mathcal{D}(\mu_{B+i},\kappa_{B+i})\right)$
@@ 5.1 Experimental Setup
Datasets.
We conduct experiments on two standard benchmarks: CIFAR-10 (Krizhevsky, Hinton et al. 2009) and ImageNet-100, a 100-class subset of ImageNet (Deng et al. 2009). These datasets allow us to evaluate the proposed method on both small- and large-scale visual recognition tasks.
We investigated transfer learning performance on the Food-101 dataset (Bossard, Guillaumin, and Van Gool 2014), CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), FGVC Aircraft (Maji et al. 2013), the Describable Textures Dataset (DTD) (Cimpoi et al. 2014) and Oxford-IIIT Pets (Parkhi et al. 2012).
We adopt CIFAR-10-C (Hendrycks and Dietterich 2019) and ImageNet-100-C as corruption benchmarks to evaluate the robustness of the learned representations. The ImageNet-100-C dataset is constructed by selecting the 100 classes used in ImageNet-100 from the original ImageNet-C benchmark. For both datasets, we use only the corrupted samples with severity level 1 for evaluation.
Model Architectures.
We follow the MoCo v2 (Chen et al. 2020b) framework, using ResNet-18 as the encoder for CIFAR-10, with a 128-dimensional projection head. For ImageNet-100, we adopt MoCo v3 (Chen, Xie, and He 2021) with a ViT-S/16 backbone to evaluate scalability across both CNN and transformer architectures.
Data Augmentation.
For data augmentation, we apply a combination of RandAugment (Cubuk et al. 2020), ColorJitter, Random Grayscale, Gaussian Blur, and Horizontal Flip, following common practice in CL setups.
@@ 5.2 Evaluations
We evaluate the model using two standard protocols: k-Nearest Neighbors (kNN) and Linear Evaluation. The kNN evaluation measures similarity preservation in the embedding space, while linear evaluation assesses the representational quality via a trained classifier.
@@
Using $k=200$, we conduct kNN evaluation on CIFAR-10 and ImageNet-100 with features from the trained encoder. DSF method achieves 79.00% on ImageNet-100 and 90.04% on CIFAR-10, outperforming MoCo v3 baselines (74.08% and 88.44%, respectively) and other multi-view methods. See Table 1.
@@
We freeze the pretrained encoder and train a linear classifier on top. DSF achieves 85.80% on ImageNet-100 and 91.21% on CIFAR-10, surpassing MoCo v3 results of 83.42% and 88.10% and also other multi-view methods. These results indicate that DSF produces more discriminative and transferable representations. See Table 1.
Transfer Learning.
Transfer learning is one of the ultimate goals of SSL. We evaluate the transfer performance on five natural image datasets under the fine-tuning setting, where all model weights are updated without freezing. More details are in Appendix D.3
Robustness under Distribution Shift.
To assess the robustness of the learned representations under distribution shift, we evaluate the models on common corruption benchmarks: CIFAR-10-C and ImageNet-100-C, following the protocol of (Hendrycks and Dietterich 2019). We use the mean Corruption Error (mCE) as the evaluation metric, measured using the model parameters obtained from linear evaluation on CIFAR-10 and ImageNet-100, respectively. No additional fine-tuning is performed during this evaluation.
As shown in Table 1, DSF consistently outperforms both the baseline and other multi-view contrastive learning methods, achieving an mCE of 16.20% on CIFAR-10-C and 33.15% on ImageNet-100-C. These results indicate that the joint structure–capturing capability of DSF also enhances the model’s robustness to input corruptions.
(a) kNN Top-1 Accuracy with respect to the number of view.
(b) Linear Evaluation Top-1 Accuracy with respect to the number of view.
@@
Model Method Dataset (Top-1 Accuracy %)
CIFAR10 CIFAR100 Pet Aircraft DTD Average
MoCo v3 - 96.36 80.80 84.11 47.58 59.68 73.70
MoCo v3 Loss_avg 96.76 81.55 83.11 47.17 58.67 73.45
MoCo v3 Feature_avg 96.72 81.06 82.80 46.45 57.66 72.93
MoCo v3 DSF 97.11 81.76 84.17 49.49 60.59 74.62
Table 3: Comparison of transfer learning performance. Top-1 accuracy (%) on five downstream datasets using MoCo v3 representations pre-trained on ImageNet-100. Bold indicates the best performance, and underline indicates the second best.
@@ 5.3 Same Resource Comparison
DSF samples multiple views per image, which may increase computational cost. To ensure fair comparison, we fix the product of batch size and number of views ($B\times M$) across all methods, maintaining equal GPU memory usage. For example, in ImageNet-100 experiments, MoCo v3 uses $512\times 2$, while DSF and other multi-view baselines use $128\times 8$, preserving memory usage. In CIFAR-10 experiments, MoCo uses $1024\times 2$, while DSF and other multi-view baselines adopt $256\times 8$.
To align training time, we reduce epochs for multi-view methods proportionally. This keeps runtime differences within 1 hour for ImageNet-100 (under 7%) and 30 minutes for CIFAR-10 (under 10%), ensuring that performance gains arise from DSF itself rather than extra computation.